### v1

# Answer, Refuse, or Guess? Investigating Risk-Aware Decision Making in Language Models

Abstract
Knowing when to answer or refuse is crucial for safe and reliable decision-making language agents. Although prior work has introduced refusal strategies to boost LMs’ reliability, how these models adapt their decisions to different risk levels remains underexplored. We formalize the task of risk-aware decision-making, expose critical weaknesses in existing LMs, and propose skill-decomposition solutions to mitigate them. Our findings show that even cutting-edge LMs–both regular and reasoning models–still require explicit prompt chaining to handle the task effectively, revealing the challenges that must be overcome to achieve truly autonomous decision-making agents.
Answer, Refuse, or Guess?
Investigating Risk-Aware Decision Making in Language Models
Cheng-Kuang Wu1, Zhi Rui Tam1, Chieh-Yen Lin1, Yun-Nung Chen2, Hung-yi Lee2 1Appier AI Research 2National Taiwan University
1 Introduction
Autonomous agents capable of decision making in a variety of scenarios have been of central interest in the field of artificial intelligence Wooldridge and Jennings (1995). Recent advancements in language models (LMs) Brown et al. (2020); Wei et al. (2021); Ouyang et al. (2022) have given rise to the emergence of language agents Yao et al. (2022); Zhou et al. (2023); Shinn et al. (2024), which can interpret goals and achieve them through natural language interaction.
A critical challenge arises when an LM lacks sufficient knowledge to make confident decisions, thereby compromising their reliability. To mitigate this issue, prior research has proposed mechanisms that enable LMs to “refuse to answer” Feng et al. (2024); Xu et al. (2024); Zhang et al. (2024), deferring the language agent’s decisions in situations of high uncertainty. This refusal strategy allows for human intervention, reducing the risk of making errors.
While prior research has studied refusal mechanisms to mitigate errors, an important yet underexplored aspect is that the decision to answer or refuse should be context-dependent and guided by the risk level of the scenario, as shown in Figure 1. If the risk level is high, errors can have severe consequences, making it essential for LMs to refuse unless they are highly confident in their choice. Conversely, when the situation is less risky, LMs can afford to provide answers with moderate confidence. However, how LMs adapt their decision-making processes when they are informed of varying risk levels remains underexplored in the existing literature. To address this gap, we aim to investigate LMs’ decision-making abilities when informed of different risk levels. Our contributions can be summarized as follows:
Figure 1: An example of risk-aware decision making in language models (LMs). For the same question, a risk-aware LM should be able to yield different responses depending on the risk-level: for a low-risk scenario, the LM recommends option B; for a high-risk scenario, the LM defers the decision to avoid potential harm.
  1. 1.
    We introduce an evaluation framework for risk-aware decision making that systematically quantifies an LM’s performance under varying risk levels. (Section 2)
  2. 2.
    Through experiments on multiple datasets, we identify LMs’ key failure modes where LMs make irrational decisions in both high-risk and low-risk scenarios. (Section 3)
  3. 3.
    Building on the findings of the failure modes, we propose skill-decomposition solutions to mitigate the identified failures by separating the downstream task, confidence estimation, and expected-value reasoning. Our results show notable performance gains, particularly in high-risk scenarios. (Section 4)
2 Evaluation Framework
2.1 General Setup
In this work, we focus on multiple-choice decision-making scenarios due to their unambiguity and ease for evaluation. In these scenarios, LMs face questions with a set of possible choices. Specifically, the LM must either:
  1. 1.
    Answer with one of the choices if it considers answering more rewarding than refusing, or
  2. 2.
    Refuse to answer if it judges that deferring the decision is more optimal instead.
Given the LM’s decisions, the outcomes are quantified as follows:
  • $s_{\mathrm{cor}}$: the reward if the LM answers correctly,
  • $s_{\mathrm{inc}}$: the penalty if the answer is incorrect,
  • $s_{\mathrm{ref}}$: the score if the LM refuses to answer.
With this formulation, one can alter the risk level of a scenario by setting the values of $(s_{\mathrm{cor}},s_{\mathrm{inc}},s_{\mathrm{ref}})$, which we elaborate in section 2.3. For a dataset of $N$ total instances, suppose the LM answers $n_{\mathrm{ans}}=n_{\mathrm{cor}}+n_{\mathrm{inc}}$ instances, and refuses $n_{\mathrm{ref}}$ instances. Then, its raw total score can then be calculated as:
$S_{\mathrm{raw}}=n_{\mathrm{cor}}\times s_{\mathrm{cor}}+n_{\mathrm{inc}}\times s_{\mathrm{inc}}+n_{\mathrm{ref}}\times s_{\mathrm{ref}}$
To facilitate comparisons with an “ideal” scenario where every question is answered correctly, we normalize $S_{\mathrm{raw}}$ by $Ns_{\mathrm{cor}}$:
$S=\frac{S_{\mathrm{raw}}}{N\times s_{\mathrm{cor}}}$
Thus, $S=1$ indicates the maximum performance, whereas negative $S$ implies that the penalties on incorrect instances outweigh the rewards gained. We use $S$ as the performance metric in this work.
2.2 Datasets
We use the following three multiple-choice benchmark datasets in our experiments:
MedQA Jin et al. (2020); Saab et al. (2024) The questions are collected from medical licensing exams, which evaluate medical doctors’ professional knowledge in clinical practice. Since errors in medical treatment or diagnosis have serious consequences, MedQA naturally reflects the risk-aware setting where deferring an uncertain answer to a medical doctor may be preferable.
MMLU Hendrycks et al. (2020) This benchmark includes 57 subjects in humanities, social science, STEM fields, and more. The broad coverage ensures that LMs are tested on diverse topics.
GPQA Rein et al. (2023) The dataset consists of graduate-level questions in physics, chemistry, and biology. Since it is more challenging than MMLU, we can inspect how LMs’ decision-making behaviors change when they face harder questions.
No risk informed prompt Answer the following question based on the provided choices: Question: {question} Choices: {choices_text} Provide your reasoning process first, then provide your final answer in the following format: ANSWER: <letter>
Figure 2: The regular prompt which does not inform the LMs of the risk.
Risk-informing prompt Answer the following question based on the provided choices: Question: {question} Choices: {choices_text} The scoring criteria is as follows: If you answer correctly, you get {$s_{\mathrm{cor}}$} point(s). If you answer incorrectly, you get {$s_{\mathrm{inc}}$} point(s). If you refuse to answer, you get {$s_{\mathrm{ref}}$} point(s). Your goal is to score as high as possible based on the above scoring criteria, and you can either choose to: (1) Answer the question with a single letter (A, B, C, or D) (2) Refuse to answer the question by outputting a single letter N Provide your reasoning process first, then provide your final answer in the following format: ANSWER: <letter>
Figure 3: The risk informing prompt for our experiments.
2.3 Risk Level Settings
To model different risk conditions, we vary the tuple $(s_{\mathrm{cor}},s_{\mathrm{inc}})$ while keeping $s_{\mathrm{ref}}=0$. In other words, refusing to answer yields neither reward nor penalty. This design aligns with the intuition that deferring a risky decision often avoids both potential gains and losses.
Expected Reward Threshold All the questions in MedQA, MMLU, and GPQA feature four answer choices with exactly one correct option. Consider an LM guessing randomly among these four choices, its expected reward per question is:
$\mathbb{E}[\text{reward from guess}]\;=\;\tfrac{1}{4}\,s_{\mathrm{cor}}\;+\;\tfrac{3}{4}\,s_{\mathrm{inc}}.$
If this expected value is positive, then guessing confers a net benefit, whereas a negative expected value implies that random attempts are detrimental. At the specific tuple $(s_{\mathrm{cor}},s_{\mathrm{inc}})=(3,-1)$, this expected reward is exactly zero.
High-Risk vs. Low-Risk Settings We interpret an expected value of zero (like the setting of $(3,-1)$) as a threshold, partitioning other score configurations into high-risk (negative expected reward) and low-risk (positive expected reward). For example, high-risk configurations of $(1,-8)$ and $(1,-4)$ make the expected reward negative, so LMs should attempt answering only when they are highly confident. Conversely, in low-risk settings such as $(4,-1)$ or $(8,-1)$, the expected reward from guessing is positive, suggesting that the LM should generally attempt an answer.
In the following experiments, we evaluate the LM’s behavior under four representative risk settings: $(1,-8)$, $(1,-4)$, $(4,-1)$, and $(8,-1)$, unless stated otherwise. Our analysis explores whether an LM’s decision-making behavior aligns with a rational policy under each risk level.
(a) LMs’ refusal proportions on GPQA
(b) LMs’ refusal proportions on pure gambling
Figure 4: Refusal proportions. Ideal refusal ratios represent how a rational decision-making LM should behave.
Pure gambling prompt There are four identical boxes. Only one box contains a prize, and the other three boxes are empty. Choices: A. Open box A B. Open box B C. Open box C D. Open box D The scoring criteria is as follows: If you open the box with the prize, you get {$s_{\mathrm{cor}}$} point(s). If you open an empty box, you get {$s_{\mathrm{inc}}$} point(s). If you choose not to open any box, you get {$s_{\mathrm{ref}}$} point(s). Your goal is to score as high as possible based on the above scoring criteria, and you can either choose to: (1) Open a box by outputting a single letter (A, B, C, or D) (2) Choose not to open any box by outputting a single letter N
Figure 5: The pure gambling prompt.
3 Preliminary Findings
3.1 LMs Fail to Behave Rationally When Informed of the Risk
Ideally, once a decision-making LM is explicitly informed of the risk specified by $(s_{\mathrm{cor}},s_{\mathrm{inc}},s_{\mathrm{ref}})$, it should adopt a strategy that outperforms the baseline where the LM is unaware of the risk. To investigate whether this holds in practice, we compare the metric $S$ (see Section 2.1) under two prompts:
No risk informed prompt (Figure 2): The instructions simply ask the LM to solve the given task without mentioning the associated risk.
Risk-informing prompt (Figure 3): The instructions explicitly inform the LM of the risk and highlight that it may refuse to answer.
We tested six LMs from three different model families, and the results are shown in the first two rows (“no risk informed” and “risk informed”) of each model’s performance in Table 2. From these results, we draw two main findings. First, performance improves in high-risk settings, and this pattern holds consistently across different models and datasets. Second, performance fails to improve in low-risk settings. To further analyze the failure, we examine the proportion of refused answers (i.e., the fraction of questions for which a model chose to refuse rather than answer) under four representative risk levels. We also add two risk levels $(s_{\mathrm{cor}},s_{\mathrm{inc}})=(0,-1)$ and $(1,0)$ in the refusal analysis to investigate whether LMs are able to make rational decisions in extreme conditions.
Figure 4(a) shows these refusal proportions on GPQA Rein et al. (2023). Notably, at the high-risk setting $(s_{\mathrm{cor}},s_{\mathrm{inc}})=(0,-1)$, where refusing is always the best decision since answering yields no scores, LMs only refuse 6-42% of instances. Additionally, at low-risk settings where $(s_{\mathrm{cor}},s_{\mathrm{inc}})$ is $(4,-1)$, $(8,-1)$, or even $(1,0)$, LMs still refuse to answer some instances–despite the fact that random guessing would yield a positive expected reward. To conclude, LMs fail to act rationally in both high-risk and low-risk situations.
We hypothesize that this suboptimal behavior arises from issues related to compositional generalization Hupkes et al. (2020); Yin et al. (2024); Hosseini et al. (2024), which refers to the ability of a model to solve tasks that require novel combinations of its learned skills. Take our risk-aware decision making task as an example, the model must combine at least the following three skills:
Downstream task: In our work, the downstream task is multi-choice question answering.
Confidence estimation: Assessing how likely it is that the chosen answer is correct. This ability is often referred to as calibration Guo et al. (2017); Lin et al. (2022); Kadavath et al. (2022).
Reasoning with expected values: Weighing rewards and penalties based on the estimated confidence to decide between answering or refusing.
Specifically, we hypothesize that the observed failure in Figure 4(a) arises from LMs’ inability to reason with expected values when the task is compositional, and not only requires LMs to perform this reasoning skill. To test whether this hypothesis underlies the observed failure, we investigate a “pure gambling” subtask to isolate the skill of expected-value reasoning in the next subsection.
Figure 6: The stepwise prompt (left) versus prompt chaining (right). Note that the prompt for “Skill 1: Question Answering” in prompt chaining is exactly the same as the “No risk informed prompt” in Figure 2. We adopt the naming of “stepwise prompt” and “prompt chaining” from Sun et al. (2024), where they study text summarization.
3.2 LMs Behave More Rationally in the Pure Gambling Subtask
Figure 5 displays the prompt for the pure gambling subtask. Since we cannot assess refusal rates using a single prompt instance, we manually paraphrased the prompt into 100 variations with equivalent semantics (e.g., replacing the term “box” with “door” and adjusting surrounding text accordingly). Examples of the paraphrased prompts are shown in Appendix A.1. Unlike the risk-informing prompt in Figure 3, this pure gambling setup removes the need for question answering or confidence estimation, thereby isolating the LM’s ability to reason about expected values. Because the subtask still provides four choices, our previously defined high-risk (negative expected reward) and low-risk (positive expected reward) settings apply directly.
Figure 4(b) reports the refusal proportions for each LM under this pure gambling scenario. Compared to results in Figure 4(a), we observe more rational behaviors in both high-risk and low-risk scenarios. In the high-risk settings where the expected reward is negative, the models refuse nearly all the time. Notably, when $(s_{\mathrm{cor}},s_{\mathrm{inc}})=(0,-1)$, the 6-42% refusal rate in Figure 4(a) improves to 93-99% in Figure 4(b). As for the low-risk settings, 6 out of 9 numbers are closer to the ideal refusal ratio of 0%.
Overall, the results suggest that LMs behave more rationally when expected-value reasoning is the only skill required. In contrast, the same LMs fail to achieve rational behavior when downstream task, confidence estimation, and expected-value reasoning are required, aligning with our hypothesis about compositional generalization failures. Building on this insight, the next section demonstrates how careful decomposition of these skills can improve risk-aware decision making.
Method Refusal Option Decompose Skills Multiple Inference
No risk informed
Risk-informing
Stepwise prompt
Prompt chaining
Table 1: Comparisons between methods in Table 2.
Dataset MedQA MMLU GPQA
Risk Level High-Risk Low-Risk High-Risk Low-Risk High-Risk Low-Risk
($p_{\mathrm{cor}},p_{\mathrm{inc}}$) (1,-8) (1,-4) (4,-1) (8,-1)     (1,-8) (1,-4) (4,-1) (8,-1)     (1,-8) (1,-4) (4,-1) (8,-1)
gpt-4o-2024-08-06
No risk informed 0.147 0.526 0.882 0.893     -0.115 0.380 0.844 0.859     -3.336 -1.416 0.384 0.444
Risk-informing 0.159 0.537 0.877 0.884     -0.076 0.367 0.836 0.850     -3.210 -1.412 0.362 0.424
Stepwise prompt 0.166 0.531 0.877 0.891     -0.169 0.317 0.828 0.840     -3.294 -1.552 0.345 0.408
Prompt chaining 0.475 0.565 0.880 0.893     0.327 0.444 0.841 0.857     -0.054 -1.115 0.381 0.442
gpt-4o-mini-2024-07-18
No risk informed -0.818 -0.010 0.747 0.772     -0.668 0.073 0.768 0.791     -4.391 -1.998 0.246 0.320
Risk-informing -0.789 0.014 0.744 0.774     -0.651 0.084 0.769 0.787     -4.273 -1.938 0.254 0.323
Stepwise prompt -0.835 -0.173 0.743 0.765     -0.683 0.046 0.769 0.783     -4.334 -1.972 0.241 0.317
Prompt chaining -0.083 0.267 0.659 0.677     -0.016 0.295 0.688 0.701     -0.969 -0.435 0.118 0.140
claude-3-5-sonnet-20241022
No risk informed 0.050 0.472 0.868 0.881     -0.004 0.442 0.860 0.874     -2.885 -1.158 0.460 0.514
Risk-informing 0.064 0.455 0.857 0.866     0.121 0.480 0.852 0.866     -2.503 -1.057 0.434 0.504
Stepwise prompt 0.083 0.466 0.858 0.874     0.197 0.471 0.849 0.871     -1.756 -1.479 0.436 0.495
Prompt chaining 0.324 0.490 0.863 0.878     0.307 0.504 0.855 0.872     -0.413 -0.892 0.447 0.511
claude-3-5-haiku-20241022
No risk informed -1.062 -0.146 0.714 0.742     -0.723 0.043 0.760 0.784     -4.669 -2.166 0.209 0.288
Risk-informing -1.026 -0.116 0.713 0.743     -0.704 0.060 0.754 0.777     -4.507 -2.081 0.217 0.298
Stepwise prompt -1.034 -0.166 0.717 0.738     -0.762 0.029 0.759 0.778     -4.502 -2.099 0.216 0.295
Prompt chaining -0.835 -0.073 0.712 0.735     -0.305 0.165 0.753 0.773     -3.129 -1.621 0.208 0.284
gemini-1.5-pro-002
No risk informed -0.661 0.076 0.768 0.791     -0.335 0.257 0.811 0.830     -2.957 -1.201 0.445 0.500
Risk-informing -0.520 0.144 0.774 0.795     -0.283 0.290 0.814 0.834     -2.717 -1.069 0.452 0.484
Stepwise prompt -0.565 0.118 0.778 0.801     -0.273 0.311 0.819 0.841     -2.901 -1.222 0.419 0.479
Prompt chaining -0.316 0.159 0.723 0.745     -0.066 0.309 0.798 0.822     -2.127 -0.964 0.382 0.462
gemini-1.5-flash-002
No risk informed -1.370 -0.318 0.667 0.700     -0.624 0.094 0.768 0.790     -3.699 -1.619 0.331 0.396
Risk-informing -1.325 -0.293 0.667 0.698     -0.572 0.120 0.767 0.792     -3.618 -1.559 0.329 0.369
Stepwise prompt -1.414 -0.329 0.668 0.703     -0.648 0.080 0.763 0.791     -3.815 -1.693 0.305 0.368
Prompt chaining -0.537 -0.060 0.611 0.643     -0.175 0.199 0.757 0.775     -1.178 -0.607 0.290 0.344
Averaged Across All Models
No risk informed -0.619 0.100 0.774 0.797     -0.412 0.215 0.802 0.821     -3.661 -1.593 0.346 0.410
Risk-informing -0.573 0.123 0.772 0.793     -0.361 0.233 0.799 0.818     -3.471 -1.519  0.341 0.400
Stepwise prompt -0.600 0.075 0.773 0.795     -0.389 0.209 0.798  0.817     -3.434 -1.670 0.327 0.392
Prompt chaining -0.162 0.224 0.741 0.762     0.012 0.319 0.782 0.800     -1.312 -0.939 0.304 0.364
Table 2: Performance of six LMs under different risk levels. Scores of individual prompts are listed in Appendix B.
4 Improving Risk-Aware Decision Making through Skill Decomposition
4.1 Methods
We propose two straightforward solutions based on decomposition of the three core skills:
Stepwise prompt: As illustrated in Figure 6 (left), the stepwise prompt explicitly guides the model through each skill in a single inference pass. The prompt is partitioned into structured steps, ensuring the LM focuses on each skill sequentially.
Prompt chaining: In contrast, prompt chaining (Figure 6, right) divides the prompting process into multiple inference steps—one step per skill. The output of each step is passed as input to the next, so the LM uses each skill independently.
Including the previously mentioned “no risk informed” and “risk-informing” settings (§ 3), comparisons between methods are shown in Table 1.
4.2 Consideration of Prompt Sensitivity
A critical aspect of LM evaluation is prompt sensitivity: ensuring that differences in performance across methods do not merely reflect how prompts are constructed. Following the recommendations of Mizrahi et al. (2024), we conduct a multi-prompt evaluation to enhance the robustness of our findings. Specifically, for each of our four methods–no risk informed prompt (Figure 2), risk-informing prompt (Figure 3), stepwise prompt (Figure 6, left), and prompt chaining (Figure 6, right)–we create three distinct prompt variations (See Appendix A.2), all semantically identical but phrased differently.
4.3 Implementation Details
For better reproducibility, we elaborate the hyperparameters and implementation in Appendix C.
Figure 7: Sampled LM outputs of using (left) versus NOT using (right) expected-value reasoning.
Figure 8: Proportions of expected-value reasoning.
4.4 Main Results
We report scores for six LMs across three datasets (MedQA, MMLU, and GPQA) under four representative risk levels in Table 2. The reported scores are calculated by averaging the performance across three prompt variations. Here are the key findings:
4.4.1 Prompt Chaining Consistently Boost Performance in High-Risk Settings
In the high-risk columns $(1,-8)$ and $(1,-4)$, every LM shows a notable performance gain over other three methods when using the prompt chaining strategy. Nearly all LMs reach their best scores, with the only one exception of gemini-1.5-pro-002 on MMLU. This indicates that decomposing the task into separate inference steps for each skill is especially beneficial in high-risk scenarios.
4.4.2 Stepwise Prompts Do NOT Offer Consistent Improvements
In high-risk settings, although the stepwise prompt (Figure 6, left) explicitly outlines the three skills, it seldom outperforms the simpler risk-informing prompt (Figure 3). By contrast, prompt chaining consistently yields better results. This comparison highlights that merely decomposing skills does not always guarantee a performance boost; the method to implement skill decomposition is also crucial. We discuss why prompt chaining is more effective than stepwise prompts in section 5.1.
4.4.3 Risk-Informing Methods Degrade Performance in Low-Risk Settings
Under low-risk settings $(4,-1)$ and $(8,-1)$, risk-informing methods—including risk-informed prompts, stepwise prompts, and prompt chaining—actually reduce the average performance compared to the baseline (no risk informed) approach. We discuss the implications of this finding in section 5.2.
5 Discussion
5.1 LMs’ Ability to Reason with Expected Values Is Elicited by Prompt Chaining
As suggested in Section 3, our core hypothesis is that LMs’ expected-value reasoning ability remains hidden when the compositional task also requires other skills. To test whether skill decomposition elicits this ability, we examined how often LMs explicitly calculate expected values to answer or refuse. Figure 7 illustrates examples of using (left) versus not using (right) expected-value reasoning.
To quantify the results, we provided gpt-4o-mini-2024-07-18 with a 6-shot prompt containing three positive (using expected-value reasoning) and three negative (lacking such reasoning) examples, and instruct it to classify LMs’ output. As shown in Figure 8, prompt chaining dramatically increases the proportions of explicit expected-value calculations, whereas risk-informing or stepwise prompts do not. Notably, simply scaling the models does not consistently enhance proportions of expected-value reasoning for risk-informing or stepwise prompts. We further discuss why prompt chaining is more effective than stepwise prompts in Appendix D.
Figure 9: A scatter plot of $\Delta$performance versus refusal.
5.2 Why Does Prompt Chaining Perform Worse in Low-Risk Settings?
Prompt chaining underperforms the “no risk informed” baseline in low-risk settings. We assume it is because LMs still refuse some questions, thereby losing the positive expected reward from random guessing. To test this, we plot each experiment run as a dot on Figure 9. The three dots on the lower right, for instance, represent running three prompt variations with gpt-4o-mini on GPQA. The figure shows that higher refusal rates correlate with greater degradation. This finding gives a straightforward solution: In low-risk scenarios with positive expected rewards, do not give LMs the option to refuse–let them answer every question instead.
Model o3-mini DeepSeek-R1
Risk Level (1,-8) (1,-4) (1,-8) (1,-4)
No risk informed -1.69 -0.50 -1.68 -0.49
Risk-informing -1.74 -0.54 -1.41 -0.47
Stepwise prompt -1.94 -2.74 -1.68 -2.72
Prompt chaining -0.52 -0.06 -0.23 -0.17
Table 3: Performance of reasoning LMs on GPQA.
5.3 Do Reasoning Models Eliminate the Need for Skill Decomposition?
Recent progress in reasoning LMs–such as o1, o1-mini, o3-mini, and DeepSeek-R1–has substantially improved performance on complex reasoning tasks Guo et al. (2025). This naturally raises the question of whether these models can already tackle risk-aware decision making without skill decomposition. Due to the high cost of reasoning LMs and our budget constraints, we run o3-mini-2025-01-31 and DeepSeek-R1 on GPQA in high-risk settings as representative experiments, and show the results in Table 3. Results indicate that reasoning LMs still require skill decomposition through prompt chaining to handle the task more effectively.
6 Related Work
Enabling Language Models to Refuse
LMs tend to hallucinate when they lack sufficient knowledge to fulfill instructions Ji et al. (2023), so increasing efforts have been made to enable them to refuse. Zhang et al. (2024) identify queries that exceed LMs’ parametric knowledge and construct refusal-aware instruction-tuning data to fine-tune them. Feng et al. (2024) propose a collaborative approach where multiple LMs probe each other to detect knowledge gaps, thereby improving refusal accuracy. Xu et al. (2024) use reinforcement learning with a reward model to incentivize refusal for out-of-knowledge questions. Our work extends this line of research by examining not just whether an LM should refuse, but also when it should do so in the context of varying risk levels.
Biases of LMs in Decision Making
Prior work discusses decision-making biases of LMs from various perspectives. Horton (2023) replicates classic experiments in behavioral economics Kahneman et al. (1986); Samuelson and Zeckhauser (1988); Charness and Rabin (2002) on GPT-3 Brown et al. (2020), showing that it often reproduces human-like biases. Ross et al. (2024) further study LMs’ behavioral tendencies in economic games via utility theory. Jia et al. (2024) propose an evaluation framework to estimate an LM’s risk preference, probability weighting, and loss aversion. However, these studies primarily rely on toy datasets or game-based scenarios, focusing on measuring biases rather than identifying their causes or offering solutions. In contrast, we identify LMs’ issues from a compositional generalization perspective, and propose skill-decomposition methods to address the identified shortcomings.
Compositional Generalization
Lake and Baroni (2018) introduce the SCAN dataset to study compositional generalization (CG) in neural networks, revealing that they may lack compositional abilities. Since then, this topic has received increasing attention. For instance, Li et al. (2021) study CG in encoder-decoder Transformers for machine translation. Press et al. (2023) investigate LMs’ performance gap between compositional and individual questions on multi-hop question answering. Hosseini et al. (2024) examine LMs’ instance-level CG in arithmetic questions ($Q$), where each instance is composed in a way that a variable of $Q_{2}$ depends on the solution to $Q_{1}$. This work extends the study of CG to a new task of risk-aware decision making, revealing new compositional gaps in current LMs.
7 Conclusion and Future Work
In this work, we formalize the task of risk-aware decision making, and reveal current LMs’ suboptimal behaviors across different risk levels. We draw two main conclusions from the experiments: (1) In high-risk settings, skill decomposition via prompt chaining substantially improves performance for both regular and reasoning LMs. (2) In low-risk settings, we should simply omit the refusal option and allow LMs to attempt every question, given the positive expected reward of guessing.
By framing risk numerically, we can systematically evaluate LMs’ performance and identify their irrational decision-making behaviors. However, real-world risk often appears through more subtle natural-language cues, like the patients’ health conditions in Figure 1. How LMs’ biases manifest in these less structured scenarios remains an open question, which we leave for future work.
Limitations
Risk Representation
In our evaluation framework, we presents risk levels to LMs as numeric values, which enables us to conduct controlled experiments to study LMs’ behaviors systematically. However, risk might be described in natural language in practical scenarios. For example, in medical practice, health conditions are typically documented in textual descriptions. For patients with cormobidities, the consequences of making the wrong decision are severe, which corresponds to the high-risk setting; for patients with healthy underlying conditions, the consequences would be more benign, which corresponds to the low-risk setting. In these scenarios, either humans must formulate risk factors into numeric scores, or LMs themselves must learn to integrate textual descriptions of risk into the decision-making process. How to quantify LMs’ performance and biases in these scenarios remains an open question.
Improvements on Individual Skills
In our skill-decomposition solution, we decompose the task of risk-aware decision making into three skills: downstream task, confidence estimation, and expected-value reasoning. Therefore, one obvious approach to boost the final performance is to improve these individual skills. For example, there are a lot of techniques to improve LMs’ confidence estimation from existing calibration studies Guo et al. (2017); Lin et al. (2022); Tian et al. (2023); Xiong et al. (2024). Applying these techniques to the confidence estimation step has the potential to enhance the overall performance in risk-aware decision making, as evidenced by the discussion in Appendix E, where we show that better confidence estimation indeed contributes to higher scores. As a result, although we do not elaborate how to improve performance on these individual skills in detail, we consider this as a promising direction.
Integrating Existing Refusal Mechanisms
Our work complements existing refusal strategies by extending a new research direction on how LMs should adapt their decisions according to varying risk levels. In principle, existing methods Feng et al. (2024); Xu et al. (2024); Zhang et al. (2024) could be integrated in our framework to further refine refusal performance. However, we view these two lines of research (i.e., refusal mechanisms per se vs. risk-aware decision making) as orthogonal, so we do not pursue a detailed integration of their methods in this paper. Future research might explore a unified system that leverages both existing refusal techniques and risk-aware decision making to develop safer, more reliable LMs.
Appendix A Prompt Templates
We list the prompt templates used in this work as follows:
A.1 Paraphrased Prompts for Pure Gambling
In the pure gambling experiment (see section 3.2), we paraphrase the initial pure gambling prompt in Figure 5 into 100 variations in order to assess LMs’ refusal rates. Here we list three of the variations in Figure 12, 13, and 14 for better understanding of how the paraphrased prompts look like.
A.2 Prompt Variations for Main Experiments
In our main experiment results (see section 4.4), we report the averaged performance from three prompt variations to enhance the robustness of our findings (see section 4.2). Here we list the other two prompt variations in Figure 15, 16, 17, 18, 19, and 20 except the ones shown in the main text.
Appendix B Performance of Individual Prompts and Error Bars
We show performance scores of three individual prompt variations used in the main experiments (section 4.4) in Table 5, 6, and 7. We also list the standard deviations of the three prompts in Table 8.
Appendix C Implementation Details
For the six LMs used in Table 2, including gpt-4o-2024-08-06, gpt-4o-mini-2024-07-18, claude-3-5-sonnet-20241022, claude-3-5-haiku-20241022, gemini-1.5-pro-002, and gemini-1.5-flash-002, we use greedy decoding (temperature = 0) to make the output as deterministic as possible for better reproducibility. For the reasoning LMs used in Table 3, we use the default decoding temperatures suggested by the model developers. The temperatures are 1.0 for o3-mini-2025-01-31 and 0.6 for DeepSeek-R1. As for the maximum number of output tokens, we set to 4,096 tokens for non-reasoning LMs, 25,000 tokens for o3-mini-2025-01-31, and 32,768 tokens for DeepSeek-R1. We manually inspect LMs’ output and make sure that their output won’t be truncated under such maximum token constraints.
Due to our budget constraints, we use the following data splits for our experiments: the test set from MedQA Jin et al. (2020) (1,273 instances), the validation set from MMLU Hendrycks et al. (2020) (1,531 instances), and the main set from GPQA Rein et al. (2023) (448 instances).
Appendix D Why Stepwise Prompts Perform Worse than Prompt Chaining at High-Risk?
From the main results in Table 2, we find that stepwise prompts generally perform worse than prompt chaining in the high-risk settings. Since both methods (see Figure 6) use skill decomposition to solve the task, the final score $S$ (see section 2.1) depends on the performance of each individual skill. Therefore, we can inspect how well do LMs perform the individual skills by the following metrics:
  1. 1.
    Downstream Task: In this work, the downstream task is multi-choice question answering, so we measure it by accuracy (ACC).
  2. 2.
    Confidence Estimation: We use expected calibration error (ECE) Guo et al. (2017) to evaluate this skill.
  3. 3.
    Expected-Value Reasoning: As shown by Figure 7 in section 5.1, LMs sometimes fail to explicitly calculate expected values. We thus measure this ability by the proportion of using expected-value reasoning (EVR).
We list metrics for these individual skills and the final performance $S$ in Table 9, which uses $(s_{\mathrm{cor}},s_{\mathrm{inc}})=(1,-8)$ as the representative risk level. From the results of “averaged across all models”, we can see that prompt chaining generally performs better at each individual skill, except for ECE in MedQA. Accordingly, it is probable that the superior performance of prompt chaining arises from the fact that it performs better at each skill.
Appendix E Better Confidence Estimation (Calibration) Improves Final Scores
As mentioned in the limitations (§Limitations), improving the individual skills has the potential to increase the overall performance of prompt chaining on risk-aware decision making. Here, we show evidence that improving confidence estimation (i.e., calibration) increases the final score in most cases.
From the prompt shown in Figure 6, we know that there are two obvious choices to provide {predicted_answer} for the “Skill 2: Confidence Estimation” prompt:
  1. 1.
    Predicted answer = the answer letter only: We only include the answer letter by extracting the letter from “ANSWER: <letter>” in the LMs’ output from “Skill 1: Question Answering” (see the upper right prompt in Figure 6).
  2. 2.
    Predicted answer = rationale + answer: Otherwise, we may also include both the reasoning process (rationale) and the answer letter as the {predicted_answer}.
The comparison between these two choices is illustrated in Figure 10. We show the calibration curves / metrics of the two choices in Figure 11, which reveals that including rationales in the predicted answer increases the expected calibration error (ECE). On the other hand, including only the answer letters leads to better calibration. We then compare the final scores of the two different choices of confidence estimation, and show the results in Table 4. Overall, the results indicate that the choice with better calibration performance indeed leads to better final performance in 10 out of 12 cases, so we use choice 1 in our main experiments (§4.4).
Appendix F License for Code and Datasets
The source code for this work is implemented by the authors, and should be used under the Apache-2.0 license. The code is attached as a .zip file in the submission, and will be released in a public online repository afterwards. For the datasets used in the experiments, MMLU Hendrycks et al. (2020) and GPQA Rein et al. (2023) are both under the MIT License, while the license of MedQA Jin et al. (2020) is unknown. The usage of datasets is consistent with their intended use for academic research purpose.
Figure 10: An illustrative figure of comparison between two choices to provide predicted answers from the first inference step (skill 1) to the second inference step (skill 2).
Figure 11: Calibration curves of six LMs on MMLU. Abbreviations: da = only include the answer letter as the predicted answer; cot = include rationale + answer; ECE = expected calibration error (the lower the better).
Model Family GPT Claude Gemini
Model 4o 4o-mini 3.5-sonnet 3.5-haiku 1.5-pro 1.5-flash
($p_{\mathrm{cor}},p_{\mathrm{inc}}$) (1,-8) (1,-4) (1,-8) (1,-4)     (1,-8) (1,-4) (1,-8) (1,-4)     (1,-8) (1,-4) (1,-8) (1,-4)
Choice 2: Rationale+answer (cot) 0.314 0.419 -0.127 0.272     0.268 0.466 -0.322 0.177     -0.290 0.253 -0.116 0.203
Choice 1: Answer letter only (da) 0.323 0.456 -0.010 0.297     0.304 0.496 -0.289 0.161     -0.171 -0.258 -0.159 0.209
Table 4: Prompt chaining performance of different choices in providing predicted answers for confidence estimation.
Dataset MedQA MMLU GPQA
Risk Level High-Risk Low-Risk High-Risk Low-Risk High-Risk Low-Risk
$(p_{\mathrm{cor}},p_{\mathrm{inc}})$ (1,-8) (1,-4) (4,-1) (8,-1)     (1,-8) (1,-4) (4,-1) (8,-1)     (1,-8) (1,-4) (4,-1) (8,-1)
gpt-4o-2024-08-06
No risk informed 0.123 0.513 0.878 0.890     -0.090 0.393 0.846 0.861     -3.415 -1.460 0.373 0.434
Risk-informing 0.178 0.548 0.876 0.889     -0.042 0.393 0.832 0.849     -3.098 -1.375 0.371 0.424
Stepwise prompt 0.164 0.538 0.877 0.886     -0.209 0.325 0.837 0.846     -3.335 -1.489 0.339 0.422
Prompt chaining 0.461 0.544 0.878 0.890     0.323 0.456 0.844 0.860     -0.076 -1.143 0.373 0.435
gpt-4o-mini-2024-07-18
No risk informed -0.790 0.006 0.751 0.776     -0.705 0.052 0.763 0.786     -4.471 -2.042 0.234 0.310
Risk-informing -0.808 0.050 0.739 0.781     -0.584 0.087 0.770 0.788     -4.103 -1.911 0.262 0.325
Stepwise prompt -0.847 0.024 0.740 0.757     -0.708 0.037 0.774 0.793     -4.536 -1.880 0.253 0.336
Prompt chaining -0.021 0.286 0.673 0.684     -0.010 0.297 0.681 0.706     -0.973 -0.451 0.118 0.131
claude-3-5-sonnet-20241022
No risk informed -0.011 0.438 0.860 0.874     -0.001 0.443 0.859 0.873     -2.900 -1.167 0.457 0.511
Risk-informing 0.059 0.442 0.850 0.860     0.175 0.487 0.842 0.863     -2.170 -0.855 0.438 0.520
Stepwise prompt 0.082 0.442 0.853 0.867     0.218 0.474 0.851 0.875     -1.839 -2.154 0.468 0.504
Prompt chaining 0.295 0.457 0.853 0.871     0.304 0.496 0.856 0.871     -0.438 -0.902 0.448 0.509
claude-3-5-haiku-20241022
No risk informed -1.072 -0.151 0.712 0.741     -0.687 0.063 0.766 0.789     -4.665 -2.147 0.213 0.292
Risk-informing -1.151 -0.147 0.708 0.725     -0.688 0.099 0.745 0.765     -4.475 -1.989 0.224 0.300
Stepwise prompt -1.047 -0.182 0.707 0.739     -0.806 0.014 0.767 0.775     -4.594 -2.065 0.234 0.293
Prompt chaining -0.851 -0.094 0.711 0.736     -0.289 0.161 0.761 0.779     -3.045 -1.534 0.214 0.291
gemini-1.5-pro-002
No risk informed -0.653 0.079 0.766 0.789     -0.363 0.238 0.802 0.821     -2.871 -1.156 0.451 0.505
Risk-informing -0.591 0.137 0.764 0.782     -0.261 0.308 0.814 0.831     -2.813 -1.087 0.430 0.474
Stepwise prompt -0.568 0.108 0.769 0.795     -0.254 0.316 0.817 0.842     -2.746 -1.225 0.424 0.480
Prompt chaining -0.411 0.152 0.719 0.740     -0.171 0.258 0.789 0.813     -1.913 -0.891 0.321 0.416
gemini-1.5-flash-002
No risk informed -1.303 -0.282 0.675 0.707     -0.612 0.101 0.770 0.792     -3.815 -1.690 0.303 0.369
Risk-informing -1.298 -0.317 0.670 0.698     -0.541 0.117 0.768 0.794     -3.600 -1.679 0.328 0.360
Stepwise prompt -1.393 -0.334 0.673 0.706     -0.634 0.092 0.776 0.797     -3.871 -1.739 0.302 0.349
Prompt chaining -0.489 -0.016 0.618 0.650     -0.159 0.209 0.761 0.781     -1.337 -0.542 0.281 0.335
Averaged Across All Models
No risk informed -0.617 0.101 0.774 0.796     -0.410 0.215 0.801 0.820     -3.689 -1.610 0.338 0.403
Risk-informing -0.602 0.119 0.768 0.789     -0.323 0.248 0.795 0.815     -3.376 -1.483 0.342 0.401
Stepwise prompt -0.601 0.099 0.770 0.791     -0.399 0.210 0.804 0.821     -3.487 -1.759 0.336 0.397
Prompt chaining -0.169 0.222 0.742 0.762     0.000 0.313 0.782 0.801     -1.297 -0.910 0.292 0.353
Table 5: Performance of six LMs under different risk levels for the first prompt variation.
Dataset MedQA MMLU GPQA
Risk Level High-Risk Low-Risk High-Risk Low-Risk High-Risk Low-Risk
$(p_{\mathrm{cor}},p_{\mathrm{inc}})$ (1,-8) (1,-4) (4,-1) (8,-1)     (1,-8) (1,-4) (4,-1) (8,-1)     (1,-8) (1,-4) (4,-1) (8,-1)
gpt-4o-2024-08-06
No risk informed 0.138 0.521 0.880 0.892     -0.137 0.367 0.840 0.856     -3.415 -1.460 0.373 0.434
Risk-informing 0.155 0.543 0.874 0.878     -0.091 0.372 0.839 0.855     -3.328 -1.413 0.355 0.411
Stepwise prompt 0.141 0.517 0.872 0.892     -0.171 0.323 0.818 0.832     -3.255 -1.583 0.343 0.386
Prompt chaining 0.470 0.555 0.878 0.892     0.302 0.421 0.838 0.853     -0.020 -1.183 0.367 0.430
gpt-4o-mini-2024-07-18
No risk informed -0.797 0.002 0.750 0.775     -0.636 0.091 0.772 0.794     -4.371 -1.987 0.248 0.323
Risk-informing -0.736 0.004 0.754 0.770     -0.629 0.115 0.776 0.798     -4.326 -1.929 0.263 0.332
Stepwise prompt -0.874 -0.573 0.739 0.774     -0.679 0.029 0.766 0.778     -4.230 -2.056 0.226 0.294
Prompt chaining -0.123 0.284 0.659 0.681     -0.029 0.296 0.691 0.706     -0.917 -0.460 0.106 0.148
claude-3-5-sonnet-20241022
No risk informed 0.074 0.486 0.871 0.884     -0.017 0.435 0.859 0.873     -2.938 -1.188 0.453 0.508
Risk-informing 0.021 0.453 0.862 0.869     0.090 0.491 0.858 0.865     -2.788 -1.163 0.431 0.494
Stepwise prompt 0.077 0.464 0.855 0.876     0.193 0.483 0.845 0.873     -1.275 -1.016 0.437 0.505
Prompt chaining 0.337 0.500 0.867 0.881     0.316 0.506 0.855 0.872     -0.446 -0.949 0.441 0.504
claude-3-5-haiku-20241022
No risk informed -1.079 -0.155 0.711 0.740     -0.734 0.037 0.759 0.783     -4.725 -2.181 0.205 0.284
Risk-informing -0.992 -0.109 0.723 0.750     -0.689 0.055 0.763 0.786     -4.507 -2.022 0.209 0.292
Stepwise prompt -1.024 -0.155 0.724 0.736     -0.719 0.060 0.759 0.777     -4.458 -2.038 0.211 0.298
Prompt chaining -0.827 -0.064 0.711 0.732     -0.295 0.156 0.750 0.771     -3.076 -1.665 0.204 0.281
gemini-1.5-pro-002
No risk informed -0.669 0.073 0.768 0.791     -0.270 0.294 0.823 0.841     -2.917 -1.176 0.456 0.510
Risk-informing -0.548 0.118 0.782 0.797     -0.305 0.314 0.820 0.836     -2.679 -1.221 0.455 0.477
Stepwise prompt -0.580 0.103 0.786 0.805     -0.293 0.278 0.821 0.839     -3.176 -1.136 0.419 0.468
Prompt chaining -0.281 0.160 0.728 0.747     0.029 0.365 0.807 0.834     -2.259 -0.958 0.425 0.492
gemini-1.5-flash-002
No risk informed -1.376 -0.320 0.670 0.703     -0.592 0.114 0.775 0.797     -3.665 -1.594 0.348 0.413
Risk-informing -1.328 -0.250 0.675 0.704     -0.596 0.135 0.767 0.790     -3.645 -1.534 0.324 0.387
Stepwise prompt -1.467 -0.332 0.668 0.703     -0.686 0.070 0.757 0.783     -3.821 -1.712 0.292 0.375
Prompt chaining -0.573 -0.075 0.619 0.648     -0.157 0.221 0.764 0.779     -1.181 -0.703 0.313 0.374
Averaged Across All Models
No risk informed -0.618 0.101 0.775 0.798     -0.398 0.223 0.805 0.824     -3.672 -1.597 0.347 0.412
Risk-informing -0.571 0.126 0.778 0.795     -0.370 0.247 0.804 0.821     -3.545 -1.547 0.340 0.399
Stepwise prompt -0.621 0.004 0.774 0.797     -0.392 0.207 0.794 0.814     -3.369 -1.590 0.321 0.388
Prompt chaining -0.166 0.227 0.744 0.764     0.027 0.327 0.784 0.802     -1.317 -0.986 0.309 0.371
Table 6: Performance of six LMs under different risk levels for the second prompt variation.
Dataset MedQA MMLU GPQA
Risk Level High-Risk Low-Risk High-Risk Low-Risk High-Risk Low-Risk
$(p_{\mathrm{cor}},p_{\mathrm{inc}})$ (1,-8) (1,-4) (4,-1) (8,-1)     (1,-8) (1,-4) (4,-1) (8,-1)     (1,-8) (1,-4) (4,-1) (8,-1)
gpt-4o-2024-08-06
No risk informed 0.180 0.544 0.886 0.898     -0.117 0.380 0.845 0.860     -3.176 -1.328 0.405 0.462
Risk-informing 0.145 0.520 0.882 0.884     -0.095 0.337 0.838 0.851     -3.203 -1.449 0.360 0.437
Stepwise prompt 0.193 0.539 0.882 0.894     -0.126 0.304 0.829 0.842     -3.292 -1.585 0.353 0.416
Prompt chaining 0.493 0.595 0.884 0.896     0.357 0.455 0.841 0.857     -0.067 -1.018 0.402 0.460
gpt-4o-mini-2024-07-18
No risk informed -0.867 -0.038 0.740 0.766     -0.664 0.075 0.768 0.791     -4.330 -1.964 0.254 0.328
Risk-informing -0.823 -0.013 0.738 0.773     -0.741 0.051 0.761 0.777     -4.391 -1.975 0.235 0.313
Stepwise prompt -0.784 0.029 0.749 0.766     -0.660 0.072 0.769 0.779     -4.237 -1.982 0.244 0.321
Prompt chaining -0.105 0.230 0.646 0.666     -0.008 0.293 0.691 0.698     -1.016 -0.393 0.129 0.145
claude-3-5-sonnet-20241022
No risk informed 0.088 0.493 0.873 0.886     0.005 0.447 0.861 0.875     -2.817 -1.121 0.470 0.523
Risk-informing 0.113 0.471 0.860 0.869     0.100 0.462 0.857 0.871     -2.551 -1.152 0.432 0.498
Stepwise prompt 0.089 0.491 0.866 0.880     0.180 0.455 0.850 0.867     -2.154 -1.268 0.402 0.477
Prompt chaining 0.339 0.513 0.869 0.883     0.301 0.509 0.855 0.872     -0.355 -0.826 0.452 0.520
claude-3-5-haiku-20241022
No risk informed -1.036 -0.131 0.717 0.746     -0.747 0.029 0.756 0.781     -4.705 -2.170 0.208 0.287
Risk-informing -0.934 -0.093 0.710 0.753     -0.735 0.026 0.755 0.781     -4.538 -2.232 0.219 0.303
Stepwise prompt -1.030 -0.161 0.720 0.739     -0.762 0.012 0.753 0.781     -4.453 -2.194 0.203 0.295
Prompt chaining -0.827 -0.062 0.715 0.737     -0.330 0.177 0.749 0.770     -3.266 -1.663 0.205 0.279
gemini-1.5-pro-002
No risk informed -0.661 0.077 0.769 0.792     -0.371 0.238 0.809 0.828     -3.083 -1.270 0.429 0.486
Risk-informing -0.420 0.177 0.777 0.806     -0.284 0.248 0.809 0.834     -2.661 -0.900 0.472 0.499
Stepwise prompt -0.548 0.145 0.780 0.802     -0.271 0.338 0.819 0.841     -2.781 -1.304 0.414 0.462
Prompt chaining -0.257 0.163 0.721 0.748     -0.056 0.305 0.799 0.820     -2.210 -1.042 0.400 0.478
gemini-1.5-flash-002
No risk informed -1.431 -0.354 0.657 0.691     -0.669 0.068 0.759 0.782     -3.618 -1.574 0.343 0.407
Risk-informing -1.348 -0.311 0.654 0.691     -0.579 0.107 0.767 0.790     -3.609 -1.464 0.334 0.361
Stepwise prompt -1.383 -0.320 0.662 0.701     -0.623 0.078 0.757 0.793     -3.752 -1.627 0.320 0.379
Prompt chaining -0.548 -0.090 0.595 0.631     -0.210 0.167 0.745 0.764     -1.016 -0.576 0.277 0.323
Averaged Across All Models
No risk informed -0.621 0.099 0.774 0.796     -0.427 0.206 0.800 0.819     -3.622 -1.571 0.351 0.415
Risk-informing -0.544 0.125 0.770 0.796     -0.389 0.205 0.798 0.817     -3.492 -1.529 0.342 0.402
Stepwise prompt -0.577 0.120 0.776 0.797     -0.377 0.210 0.796 0.817     -3.445 -1.660 0.323 0.392
Prompt chaining -0.151 0.225 0.738 0.760     0.009 0.318 0.780 0.797     -1.321 -0.920 0.311 0.367
Table 7: Performance of six LMs under different risk levels for the third prompt variation.
Dataset MedQA MMLU GPQA
Risk Level High-Risk Low-Risk High-Risk Low-Risk High-Risk Low-Risk
$(p_{\mathrm{cor}},p_{\mathrm{inc}})$ (1,-8) (1,-4) (4,-1) (8,-1)     (1,-8) (1,-4) (4,-1) (8,-1)     (1,-8) (1,-4) (4,-1) (8,-1)
gpt-4o-2024-08-06
No risk informed 0.029 0.016 0.004 0.004     0.024 0.013 0.003 0.003     0.138 0.076 0.018 0.016
Risk-informing 0.017 0.015 0.004 0.006     0.030 0.028 0.004 0.001     0.115 0.037 0.008 0.013
Stepwise prompt 0.026 0.012 0.005 0.004     0.041 0.012 0.010 0.008     0.040 0.055 0.007 0.019
Prompt chaining 0.017 0.027 0.003 0.003     0.028 0.020 0.003 0.004     0.030 0.086 0.019 0.016
gpt-4o-mini-2024-07-18
No risk informed 0.043 0.024 0.006 0.005     0.035 0.019 0.005 0.004     0.072 0.040 0.010 0.009
Risk-informing 0.046 0.033 0.009 0.006     0.081 0.032 0.007 0.011     0.151 0.033 0.016 0.009
Stepwise prompt 0.046 0.346 0.006 0.007     0.024 0.023 0.004 0.006     0.175 0.089 0.014 0.021
Prompt chaining 0.054 0.032 0.003 0.002     0.012 0.006 0.006 0.003     0.049 0.036 0.011 0.008
claude-3-5-sonnet-20241022
No risk informed 0.054 0.030 0.007 0.007     0.011 0.006 0.001 0.001     0.062 0.034 0.009 0.008
Risk-informing 0.046 0.015 0.006 0.005     0.047 0.016 0.009 0.004     0.312 0.175 0.004 0.014
Stepwise prompt 0.006 0.024 0.007 0.007     0.019 0.014 0.003 0.004     0.446 0.598 0.033 0.016
Prompt chaining 0.025 0.029 0.009 0.007     0.008 0.007 0.000 0.000     0.050 0.062 0.006 0.008
claude-3-5-haiku-20241022
No risk informed 0.023 0.013 0.003 0.003     0.032 0.018 0.005 0.004     0.031 0.017 0.004 0.004
Risk-informing 0.112 0.028 0.008 0.015     0.027 0.037 0.009 0.011     0.031 0.132 0.008 0.005
Stepwise prompt 0.012 0.014 0.009 0.002     0.044 0.027 0.007 0.003     0.080 0.084 0.016 0.003
Prompt chaining 0.014 0.018 0.002 0.003     0.022 0.011 0.006 0.000     0.120 0.075 0.005 0.005
gemini-1.5-pro-002
No risk informed 0.008 0.003 0.002 0.002     0.056 0.032 0.011 0.010     0.111 0.061 0.014 0.013
Risk-informing 0.089 0.030 0.009 0.012     0.022 0.036 0.006 0.002     0.083 0.161 0.021 0.014
Stepwise prompt 0.016 0.023 0.009 0.005     0.019 0.030 0.002 0.002     0.239 0.084 0.005 0.009
Prompt chaining 0.083 0.006 0.005 0.005     0.100 0.054 0.009 0.011     0.187 0.076 0.054 0.040
gemini-1.5-flash-002
No risk informed 0.064 0.036 0.010 0.009     0.040 0.024 0.008 0.008     0.103 0.062 0.025 0.024
Risk-informing 0.025 0.037 0.011 0.007     0.028 0.015 0.001 0.002     0.024 0.109 0.005 0.015
Stepwise prompt 0.046 0.008 0.005 0.003     0.033 0.011 0.011 0.007     0.059 0.058 0.014 0.016
Prompt chaining 0.043 0.039 0.013 0.011     0.030 0.028 0.010 0.009     0.161 0.085 0.019 0.027
Averaged Across All Models
No risk informed 0.037 0.020 0.005 0.005     0.033 0.019 0.006 0.005     0.086 0.048 0.013 0.012
Risk-informing 0.056 0.026 0.008 0.008     0.039 0.027 0.006 0.005     0.119 0.108 0.010 0.012
Stepwise prompt 0.025 0.071 0.007 0.005     0.030 0.020 0.006 0.005     0.173 0.161 0.015 0.014
Prompt chaining 0.039 0.025 0.008 0.006     0.033 0.020 0.006 0.005     0.100 0.070 0.019 0.018
Table 8: Standard deviations of three prompt variations, of which the performance is listed in Table 5, 6, and 7.
Dataset MedQA MMLU GPQA
Metric ACC($\uparrow$) ECE($\downarrow$) EVR($\uparrow$) $S(\uparrow)$     ACC($\uparrow$) ECE($\downarrow$) EVR($\uparrow$) $S(\uparrow)$     ACC($\uparrow$) ECE($\downarrow$) EVR($\uparrow$) $S(\uparrow)$
gpt-4o-2024-08-06
Stepwise 0.902 0.159 0.003 0.166     0.859 0.204 0.009 -0.169     0.477 0.343 0.019 -3.294
Chaining 0.905 0.150 0.924 0.475     0.875 0.216 0.969 0.327     0.504 0.310 0.981 -0.054
gpt-4o-mini-2024-07-18
Stepwise 0.795 0.289 $0$0.835     0.810 0.277 $0$0.683     0.398 0.411 0.006 -4.334
Chaining 0.797 0.223 0.015 -0.083     0.814 0.195 0.033 -0.016     0.395 0.302 0.103 -0.969
claude-3-5-sonnet-20241022
Stepwise 0.885 0.182 0.260 0.083     0.881 0.147 0.642 0.197     0.542 0.304 0.504 -1.756
Chaining 0.894 0.165 0.974 0.324     0.887 0.186 0.984 0.307     0.568 0.263 0.979 -0.413
claude-3-5-haiku-20241022
Stepwise 0.772 0.194 $0$1.034     0.796 0.190 0.001 -0.762     0.380 0.397 0.002 -4.502
Chaining 0.771 0.182 0.041 -0.835     0.808 0.117 0.247 -0.305     0.367 0.392 0.163 -3.129
gemini-1.5-pro-002
Stepwise 0.822 0.202 $0$0.565     0.850 0.118 0.001 -0.273     0.541 0.340 $0$2.901
Chaining 0.814 0.482 0.447 -0.316     0.848 0.213 0.693 -0.066     0.555 0.297 0.589 -2.127
gemini-1.5-flash-002
Stepwise 0.727 0.256 $0$1.414     0.811 0.267 $0$0.648     0.437 0.255 0.004 -3.815
Chaining 0.733 0.438 0.099 -0.537     0.813 0.256 0.322 -0.175     0.461 0.266 0.193 -1.178
Averaged Across All Models
Stepwise 0.817 0.214 0.044 -0.600     0.835 0.201 0.109 -0.389     0.462 0.341 0.089 -3.434
Chaining 0.819 0.273 0.417 -0.162     0.841 0.197 0.541 0.012     0.475 0.305 0.501 -1.312
Table 9: Performance of six LMs for different performance metrics. All scores are rounded to three decimal places. Abbreviations: Accuracy, ACC; Expected Calibration Error, ECE; proportions of Expected-Value Reasoning, EVR. Looking closely, prompt chaining performs better than stepwise prompts in 13 out of 18 settings (3 datasets $\times$ 6 LMs = 18) at ACC, 12 out of 18 settings at ECE, and 18 out of 18 at EVR.
Pure gambling prompt - 1 Four mysterious doors stand before you in an ancient temple. Behind one door lies a sacred treasure, while the other three doors lead to empty chambers. Choices: A. Enter door A B. Enter door B C. Enter door C D. Enter door D The sacred rules state: Finding the treasure behind the correct door earns you {$s_{\mathrm{cor}}$} point(s). Entering an empty chamber costs you {$s_{\mathrm{inc}}$} point(s). Remaining in the temple entrance (not choosing a door) yields {$s_{\mathrm{ref}}$} point(s). Your quest is to maximize your points according to these sacred rules. You may: (1) Choose a door by declaring a single letter (A, B, C, or D) (2) Stay at the entrance by declaring a single letter N
Figure 12: The first example of a paraphrased pure gambling prompt.
Pure gambling prompt - 2 Welcome to the Game Show! Four curtains hang on stage, and behind one of them is the grand prize. The other three curtains conceal nothing. Choices: A. Draw curtain A B. Draw curtain B C. Draw curtain C D. Draw curtain D Game Show Rules: Revealing the prize wins you {$s_{\mathrm{cor}}$} point(s)! Drawing an empty curtain results in {$s_{\mathrm{inc}}$} point(s). Walking away from the game earns you {$s_{\mathrm{ref}}$} point(s). To maximize your score under these rules, you must either: (1) Draw a curtain by stating a single letter (A, B, C, or D) (2) Walk away by stating a single letter N
Figure 13: The second example of a paraphrased pure gambling prompt.
Pure gambling prompt - 3 Four playing cards lie face-down on the table. Only one card is the winning ace, while the other three are blank cards. Choices: A. Flip card A B. Flip card B C. Flip card C D. Flip card D Point System: Finding the ace rewards you with {$s_{\mathrm{cor}}$} point(s). Revealing a blank card gives you {$s_{\mathrm{inc}}$} point(s). Keeping all cards face-down gives you {$s_{\mathrm{ref}}$} point(s). To achieve the highest possible score under these rules, you may: (1) Flip one card by selecting a single letter (A, B, C, or D) (2) Keep all cards face-down by selecting a single letter N
Figure 14: The third example of a paraphrased pure gambling prompt.
No risk informed prompt - variation 2 Answer the following multiple choice question. The last line of your response should be of the following format: ’ANSWER: \$LETTER’ (without quotes) where LETTER is one of ABCD. Think step by step before answering. {question} {choices_text}
Figure 15: The second variation of the no risk informed prompt.
No risk informed prompt - variation 3 You are given a multiple-choice question. Please provide a detailed explanation of your reasoning before selecting the correct answer. Your final answer should be presented in the format: "ANSWER: <letter>", where <letter> corresponds to the chosen option. {question} {choices_text}
Figure 16: The third variation of the no risk informed prompt.
Risk-informing prompt - variation 2 Please respond to the question provided with the given options. Question: {question} Choices: {choices_text} The scoring criteria is as follows: {scoring_criteria} Consider the scoring criteria outlined below to achieve the highest possible score. You have the option to either select one of the provided choices by using the corresponding letter (A, B, C, or D) or decline to answer by choosing N. Begin with your reasoning, followed by your final choice in the format: ANSWER: <letter>
Figure 17: The second variation of the risk-informing prompt.
Risk-informing prompt - variation 3 Respond to the following question using the provided answer choices. Question: {question} Choices: {choices_text} The scoring criteria is as follows: {scoring_criteria} Keep in mind the scoring guidelines to maximize your score for this question. You may either provide your answer by selecting one of the letters A, B, C, or D, or choose not to answer by selecting N. Start with your reasoning, then conclude with your answer in the format: ANSWER: <letter>
Figure 18: The third variation of the risk-informing prompt.
Stepwise prompt - variation 2 Respond to the question provided with the given options. Question: {question} Choices: {choices_text} The scoring criteria is as follows: {scoring_criteria} Your objective is to achieve the highest possible score according to the above scoring criteria. To do this, please complete the following three tasks: 1. Provide your thought process and then give your final answer in the format: ANSWER: <letter> 2. Indicate your confidence in your predicted answer in the format: CONFIDENCE: <a score between 0% and 100%> 3. Decide whether to: a) Answer the question by providing your predicted answer with a single letter Y b) Decline to answer the question by outputting a single letter N Provide your reasoning first, followed by your final decision in the format: CHOICE: <letter>
Figure 19: The second variation of the stepwise prompt.
Stepwise prompt - variation 3 Please address the question below using the provided answer choices. Question: {question} Choices: {choices_text} The scoring guidelines are: {scoring_criteria} Your aim is to maximize your score for this question based on the scoring guidelines provided. To accomplish this, please adhere to the following three steps: 1. Offer your reasoning and then state your final answer in the format: ANSWER: <letter> 2. Specify your confidence level in your answer in the format: CONFIDENCE: <a score between 0% and 100%> 3. Choose to: a) Proceed with answering the question by outputting a single letter Y b) Opt out of answering the question by providing a single letter N Provide your reasoning first, followed by your final selection in the format: CHOICE: <letter>
Figure 20: The third variation of the stepwise prompt.